Angelladiao:

How much power in horsepower is involved if 9000 j of work is done in 12 second?

Angelladiao:

Help me plsas

jimthompson5910:

power = work/time For example, if 4000 J of work is done in 8 seconds, then power = work/time power = (4000 J)/(8 seconds) power = (4000/8) joules per second power = 500 watts Then to convert this to horsepower, we divide by roughly 745.699872 Since 745.699872 watts = 1 horsepower So, 500 watts = (500/745.699872) horsepower 500 watts = 0.670511 horsepower which is approximate Follow these steps to get the answer you're looking for

jimthompson5910:

Let me know what answer you get so I can check it

Angelladiao:

Need tobe solved

jimthompson5910:

What do you get when you compute 9000/12

Angelladiao:

Given Unknow Formula Solution Answer I nedd thus

jimthompson5910:

The given is the values your teacher has provided (9000 J of work and 12 seconds)

jimthompson5910:

The formula is power = work/time

jimthompson5910:

The unknown is the power we want

Angelladiao:

What given

jimthompson5910:

The numbers your teacher gave you

Angelladiao:

Give me answer

Angelladiao:

Pls

jimthompson5910:

Sorry I can't do that. You'll have to put some effort

Angelladiao:

Verry tired

Ubaidmushtaque:

A car is moving with a non uniform velocity towards east its velocity change at different intervals calculate the instantaneous velocity at time 3 sec the distance is given by equation 2t^2-4t